MR. jhon is offered an annuity plan for $250000 that will provide him $30000 annually at start of each year @8% p.a how long it will take him to get return?


1
Expert's answer
2021-04-22T07:41:38-0400

"PV =\$250000"

"r=8%" %

"i=0.08"

"R=30000"


"250,000=30000[\frac{(1+0.08)^{-n}}{0.008}]"



"\frac{250,000\times0.08}{30,000}=1-(1+1.008)^{-n}"



"0.67=1-1.008^{-n}"


"1.008^{-n}=0.33"


"-n=\frac{log \ 0.33}{log\ 1.08}"


"-n=-14.41\\
n=14.41"

"n=14.4"